OpenStudy (anonymous):

x2-6x=27 solve for x

OpenStudy (jazzyfa30):

just divide 6 then squre your answer

OpenStudy (anonymous):

\[x^2-6x=27\] subtract the 27 from both sides \[x^2-6x - 27=0\] Factor \[x^2-6x - 27=0\] \[(x + 3)(x - 9) = 0\] Can you do it from here?

OpenStudy (anonymous):

Yes thank you!!<3